OpenStudy (anonymous):

What is the missing term,x,in the pattern -3/2, 6, x, 96, -384 A.-24 b.-16 c.16 d.24

OpenStudy (raden):

hint : next term is always times -4

OpenStudy (anonymous):

So 24? Thanks wait what if it's negative probably negative

OpenStudy (raden):

-3/2 -3/2 * -4 = 6 6 * -4 = -24 -24 * -4 = 96 96 * -4 = -384 so, x = -24 not 24

OpenStudy (anonymous):

Okay thanks:3

OpenStudy (raden):

you're welcme